# The Birth of Be Star Disks II. A High-Resolution Spectroscopic Campaign and TESS Observations of an Outburst of the Classical Be star Pavonis

### v1

@@ 3 Periodic Behavior
Thus far, our discussion of the TESS light curve taken contemporaneously with our spectroscopy has focused on the outburst in the last few days of the TESS sector 67 data (Fig. 2). The behavior seen in all currently available TESS light curves (sectors 13, 66, 67, 93, and 94) is remarkably similar. We used the program Period04 (2004IAUS..224..786L) to conduct a Fourier Analysis of the TESS data in order to quantify the photometric signals, analyzing the data in three sections (sector 13, 66+67, and 93+94) as shown in Fig. 4. We performed the Fourier analysis in this manner as the adjacent sectors (66+67 and 93+94) only include a short gap in the data collection. In all cases, the dominant feature appears at a frequency of about d-1 with an amplitude of about 20 parts per thousand (ppt). The next two strongest signals are found at d-1 (amplitude of about 6 ppt) and d-1 (amplitude of about 3 ppt). These three signals have an approximate numerical relation where the difference between the two higher-frequency signals corresponds closely to that of the lowest frequency signal (i.e. ). We used the dominant frequency, corresponding to a period of 6.13 d, to phase the five sectors of TESS photometry in Fig. 4, which is remarkable stable across the years of data. Despite this persistence, is notably lower in amplitude during sectors 66 and 67 compared to the other sectors (1 ppt less than sector 13; 5 ppt less that sectors 93 and 94), even though the amplitude appears to increase as the outburst begins at the end of sector 67. The other 2 frequencies from sectors 66 and 67 appear lower in amplitude than their respective frequencies in the other 2 sector groups, as well. Pre-whitening against the two strongest peaks reveals several additional signals at low amplitudes (third row panel of Fig. 4), generally organized in closely spaced groups, similar to the majority of other Be stars (2022AJ....163..226L).
@@
Viewing the frequency spectrum after removing all periodic variations shows a red noise profile characteristic of stochastic low-frequency variability (bottom panel of Fig. 4) as seen in virtually all luminous OB stars (e.g. 2020A&A...640A..36B). We fit the frequency spectrum with the equation
@@
This equation is based on the characterization of the Fourier properties of the Sun as described by 1985ESASP.235..199H, 2014A&A...570A..41K and others. Here, the amplitude of the Fourier transform is a function of frequency . represents the amplitude of the semi-Lorentzian fit at a frequency value of zero, represents a characteristic frequency, represents the logarithmic amplitude gradient, and represents the white noise of the data. From this fit, we derive , , , and .
@@
Included in Fig. 4 are vertical dashed lines marking three frequencies found in spectroscopic line profile variations reported by 2011A&A...533A..75L. A fourth line marks the 1.485 cycles d signal, which is a new frequency with this work. Their lowest frequency signal, , is confirmed by TESS to be genuine, and not an alias as 2011A&A...533A..75L presumed, while their is also obvious in the TESS data. There is a signal at their in two of the three sections of TESS data. While there is no hint of their 0.49 0.05 cycles d-1 in the TESS photometry, there is a signal at 1.48 cycles d-1, so that it is possible that their is a one day alias. 2011A&A...533A..75L remark that their could be the first harmonic of . The corresponding frequencies of these two signals in TESS could support this interpretation, which would imply that the physical process producing the fundamental frequency, , manifests in non-sinusoidal flux variations and non-sinusoidal line profile variations. On the other hand, given the relatively low frequency resolution of both the photometric and spectroscopic datasets of 2011A&A...533A..75L, this may simply be a numerical coincidence. At any rate, it is apparent that the photometric amplitudes of and are not tightly coupled – is not detected in the final section of TESS photometry while retains a significant amplitude. While 2011A&A...533A..75L suggested that may be a circumstellar frequency where LPVs are caused by orbiting material, the fact that this signal is detected in TESS in the absence of any emission (e.g. throughout the entirety of sectors 66+67 prior to the onset of mass ejection) suggests that at least the photometric signal at this frequency is not purely circumstellar.
@@
It should be noted that signals detected in line profile variations may not necessarily correspond to readily detectable photometric signals (and vice-versa). We also note that the frequency reported by 2011A&A...533A..75L is remarkably similar to the frequency we derive for the V/R oscillations, and seems to be marginally above the noise in the Fourier spectrum of the pre-whitened light curve for Sectors 13 and 66+67 (i.e. with the and signals from 10.1111/j.1365-2966.2006.10655.x, removed). It is somewhat difficult to see how an orbital or rotational timescale for the disk can also be present throughout entire sectors of TESS data with an outburst only being present over a few days of one of these sectors.
@@
All of the observed periodic signals in TESS are presumed to be photospheric in origin. The last 3 days of sector 67 (where the disk was starting to build up) were omitted from the frequency analysis, so that all signals from sectors 66+67 in Fig. 4 were recorded during a disk-less state. There was very weak emission present during sector 93, which had dissipated to be no longer detectable throughout sector 94. That is, no mass ejections occurred during these last two sectors. We lack any spectroscopic coverage of TESS sector 13, but there is no evidence of mass ejection in the sector 13 light curve. That is, sectors 66 and 67 have a sharp rise in flux in Fig. 4, but this is not present in the sector 13 flux, nor is a long-term decline we would assume to follow a dissipation of the disk present.
@@
4 Line Profile Variations Behavior Before, During, and After the Outburst
@@
With three strong pulsation frequencies from TESS ( = 0.163 d-1, = 1.644 d-1, = 1.485 d-1), we phased line‑profile maps to examine the behavior of Pav before, during, and after the outburst. In Fig. 1, we see a slow, d motion in the residual line profiles, but a better demonstration of the line profile variations seen in He I 4921 before, during, and after the outburst is in Fig. 5 for the three prominent signals found in the previous section.
@@
Each subplot in this figure has a combination of the TESS light curve and scaled EW values to show the outburst for He I 4921, with times indicated in gray that coincide with the spectral data used. Then, we include a residual plot for the spectral line as a function of phase calculated with the three prominent TESS frequencies. Lastly, each gray scale residual of the spectral lines is compared to the phased TESS light curve from sectors 66 and 67 when our spectra were obtained.
@@
The strongest photometric signal was d-1, with an amplitude of about 20 ppt (left column of Fig. 5). The spectroscopic signal appears clearly in each spectral line, with a general trend of an absorption excess starting on the blue side of the line and progressing to the red over the 6.13 d period. The behavior is the same in all three subsets of the data, taken before, during, and after the disk outburst. These figures do contain gaps in phase, which is not unusual for such a small frequency, but this does make the pattern within the LPVs harder to see. One thing to note in the He I and H I lines is that the pulsational behavior is within the expected values of km s-1 and km s-1. The km s-1 value is presented in 2016A&A...595A.132Z as the photospheric projected rotational velocity, and makes for a good reference when looking at line features. The km s-1 lines are included to distinguish the extent of the spectral lines, and thus the velocity range that stellar signals we expect to be confined to. During the outburst, activity and variability can be seen beyond the lines at km s-1, indicating that there must be some disk material present.
@@
The second ( d-1) and third ( d-1) strongest frequencies in TESS are more challenging to recover in our spectroscopic data, especially when analyzing sub-sections of the dataset. This is likely due to a number of factors. First, the slow d-1 signal dominates the periodic LPVs. Even in the He I lines analyzed here, the net changes in the EW from circumstellar material have a relatively high amplitude compared to more subtle pulsational LPVs. The signal d-1 is about one d-1 greater than the d-1 signal, which likely causes problems with aliasing. Finally, the fast LPVs discussed in Sec. 5.2 are also relatively strong and contribute heavily to the ‘noise’ when searching for LPV patterns at and .
@@
After sub-dividing the spectroscopic data, the only hint of is seen in the pre-outburst data for He I 4921. Although marginally detected, the LPV pattern associated with is similar to that reported in 2011A&A...533A..75L, with the alternating light/dark patterns moving across the line profile from blue to red twice per phase. Due at least in part to the issues mentioned in the previous paragraph, is not recovered in the data strings during and after the outburst, and is not apparent in any data sub-strings.
@@
Our sampling of the spectrum allowed us to obtain some very dense nights of spectroscopic data throughout the observational campaign. Thus we are also able to see if there are any short timescale variations over individual nights before, during, and after the outburst, although the spectroscopic data are not complete enough in time coverage for determination of periods. Fig. 6 details the H, He I 4713, He I 6678, and Si III 5739 lines. Of the nine nights of dense spectroscopic coverage, we see fast-moving substructures in every time-series, likely related to high-order pulsations in the star that do not have a photometric period due to vanishingly smaller amplitudes when the order gets larger. As these are seen in every time-series, there does not seem to be a correlation between these line profile variations and the disk-building or dissipation. However, especially in the photospheric Si III line, we see that there are fast-moving prograde (blue-to-red) and retrograde (red-to-blue) features that occur at the same time in this star. The width of the subfeatures seems to be line dependent.
5 Discussion
@@
5.1 Low-frequency stellar variability
@@
Besides the single outburst, the most conspicuous variability is associated with a 6.13 d period ( d-1). This signal appears in TESS photometry as a sinusoidal modulation of brightness with an amplitude of about 15 – 20 ppt (Figs. 2, 4), and is present in all available TESS sectors (spanning about six years between sector 13 and 94). This same signal is also clearly seen in spectroscopic line profile variations, as first noticed by 2011A&A...533A..75L, and as shown in the dynamical spectra in Fig. 5.
@@
This 0.163 d-1 signal roughly corresponds to the difference between two frequencies observed by TESS, at 1.644 d-1 and 1.485 d-1 (differing by 2%). Such “difference frequencies” are routinely observed in Be stars, especially with the recent proliferation of high-precision space photometry (e.g. 2016A&A...588A..56B; Baade2017; Baade2018c). In general the amplitude of difference frequencies is often found to vary significantly over time. This is not the case for Pav – the difference frequency is remarkably regular over the so-far six year TESS baseline (Fig. 4), and possibly even much longer considering the spectroscopic detection of this signal in data from 1999 to 2001 (2011A&A...533A..75L), and the photometric detection in Hipparcos data collected in the early 1990s at a frequency of 0.16179 d-1 and an amplitude of 0.016 mag (approximately 16 ppt; 2002MNRAS.331...45K).
@@
The line profile variations phased to the period associated with the difference frequency appear to move in a prograde fashion, from blue to red (Figs. 1, 5). From the stellar parameters given in 2016A&A...595A.132Z, the rotation frequency of Pav is 0.46 0.15 d-1, and the critical rotation frequency is 0.62 0.14 d-1 (Paper 1). Since the difference frequency observed at 0.163 d-1 is obviously lower than the rotation frequency, it must be retrograde (unless it is zonal, i.e. , but this seems inconsistent with the LPV pattern). The pattern that moves across the line profile is relatively simple, appearing similar to an 1 or 2 mode, as are regularly seen in Be stars (2003A&A...411..229R), but at a much lower observed frequency.
@@
Another prominent example of a Be star with a difference frequency is HD 6226. In this case, the difference frequency is not observed directly (either in photometry or line profile variations), but rather is presumed because of two closely spaced pulsation frequencies detected in photometry separated by about 0.0115 d-1 (the corresponding period is 87 d). In HD 6226, outbursts occur with an 87 d period (although some events are skipped or this is modulated on two frequencies). This suggests that the timing of the outbursts is dictated by the difference frequency. In Pav, the difference frequency is over an order of magnitude faster than in HD 6226. Further, mass ejections clearly do not occur every 6 days in Pav. The situation in Pav invokes the idea of “hierarchical clocks” as discussed in 2018pas8.conf...69B, where some unknown mechanism (possibly related to very low difference frequencies) ‘opens a valve’ to allow for mass ejection, where the difference frequency then is able to pump out material. In HD 6226, this hypothetical valve is typically open, while in Pav it is not. One interesting hypothesis that we do not have the ability to test with the current data is that the 0.163 d-1 increases in its photometric amplitude at the onset of an outburst as may be seen in Fig. 2. However, the TESS data only cover half of one period of this, so it is impossible to gauge.
@@
The LPV pattern of the slow 0.163 d-1 signal seems qualitatively different from that of the faster 1.644 d-1 signal, most clearly seen in Fig. 5 of (2011A&A...533A..75L), but also in the right panels of Fig. 5. For the 0.163 d-1 signal, the dark/light patches travel across the line from phase 0 to 1, while for the 1.644 d-1 signal the dark patch (the most evident feature in 2011A&A...533A..75L) and light patch (the most evident feature in Fig. 5 in the pre-outburst data) travel across the line profile over just half of the pulsational phase.
@@
It is worth emphasizing one major shortcoming of the TESS photometry and the spectroscopic data analyzed in this work and in 2011A&A...533A..75L. The frequency resolution, which is dictated by the time baseline, is poor. There may therefore be multiple closely spaced frequencies that we detect as only one signal, especially considering that rapid rotation can lead to very small separations between the frequencies associated with consecutive radial orders (e.g. 2017A&A...598A..74P; 2022MNRAS.511.1529S). This would also naturally lead to variable amplitudes and/or frequencies in the signals measured by TESS, as unresolved signals beat against each other.
@@
That the LPVs and photometric variations from this difference frequency are so clearly observed and are apparently stable for long periods of time makes Pav an excellent laboratory to study the non-linear mode coupling that seems responsible for the difference frequency (2024A&A...687A.265V). We are not aware of any other Be star in which the spectroscopic signature of a difference frequency is so evident. A more detailed and systematic investigation of the line profile and brightness variations associated with the 0.163 d-1 signal as well as the two likely ‘parent’ modes (at 1.644 d-1 and 1.485 d-1) may help to answer several important question: Is the pulsation geometry the same for all three observed signals? Why is the difference frequency so regular compared to the majority of Be stars? How do the observed variations map to features on the stellar surface? Given the early spectral type (B2Ve), which is relatively hot compared to the blue edge of the SPB instability strip, what mechanisms excite the two parent modes?
5.2 High-frequency stellar variability

### v2

@@ III Periodic Behavior
Thus far, our discussion of the TESS light curve taken contemporaneously with our spectroscopy has focused on the outburst in the last few days of the TESS sector 67 data (Fig. 2). The behavior seen in all currently available TESS light curves (sectors 13, 66, 67, 93, and 94) is remarkably similar. We used the program Period04 (Lenz and Breger, 2004) to conduct a Fourier Analysis of the TESS data in order to quantify the photometric signals, analyzing the data in three sections (sector 13, 66+67, and 93+94) as shown in Fig. 4. We performed the Fourier analysis in this manner as the adjacent sectors (66+67 and 93+94) only include a short gap in the data collection. In all cases, the dominant feature appears at a frequency of about d-1 with an amplitude of about 20 parts per thousand (ppt). The next two strongest signals are found at d-1 (amplitude of about 6 ppt) and d-1 (amplitude of about 3 ppt). These three signals have an approximate numerical relation where the difference between the two higher-frequency signals corresponds closely to that of the lowest frequency signal (i.e. ). We used the dominant frequency, corresponding to a period of 6.13 d, to phase the five sectors of TESS photometry in Fig. 4, which is remarkably stable across the years of data. Despite this persistence, is notably lower in amplitude during sectors 66 and 67 compared to the other sectors (1 ppt less than sector 13; 5 ppt less than sectors 93 and 94), even though the amplitude appears to increase as the outburst begins at the end of sector 67. The other 2 frequencies from sectors 66 and 67 appear lower in amplitude than their respective frequencies in the other 2 sector groups, as well. Pre-whitening against the two strongest peaks reveals several additional signals at low amplitudes (third row panel of Fig. 4), generally organized in closely spaced groups, similar to the majority of other Be stars (Labadie-Bartz et al., 2022).
@@
Viewing the frequency spectrum after removing all periodic variations shows a red noise profile characteristic of stochastic low-frequency variability (bottom panel of Fig. 4) as seen in virtually all luminous OB stars (e.g., Bowman et al., 2020). We fit the frequency spectrum with the equation
@@
This equation is based on the characterization of the Fourier properties of the Sun as described by Harvey (1985), Kallinger et al. (2014) and others. Here, the amplitude of the Fourier transform is a function of frequency . represents the amplitude of the semi-Lorentzian fit at a frequency value of zero, denotes a characteristic frequency, is the logarithmic amplitude gradient, and is the white noise of the data. From this fit, we derive ppt, d-1, , and ppt.
@@
Included in Fig. 4 are vertical dashed lines marking three frequencies found in the TESS photometry reported by Levenhagen et al. (2011, = 0.17 0.02, = 0.82 0.03, and = 1.63 0.04 cycles d-1). A fourth line marks the 1.485 cycles d-1 signal, which is a new frequency with this work. Their lowest frequency signal, , is confirmed by TESS to be genuine, and not an alias as Levenhagen et al. (2011) presumed, while their is also obvious in the TESS data. There is a signal at their in two of the three sections of TESS data. While there is no hint of their 0.49 0.05 cycles d-1 in the TESS photometry, there is a signal at 1.485 cycles d-1, so it is possible that their is a one day alias. Levenhagen et al. (2011) remark that their could be the first harmonic of . The corresponding frequencies of these two signals in TESS could support this interpretation, which would imply that the physical process producing the fundamental frequency, , manifests in non-sinusoidal flux variations and non-sinusoidal line profile variations. On the other hand, given the relatively low frequency resolution of both the photometric and spectroscopic datasets of Levenhagen et al. (2011), this may simply be a numerical coincidence. At any rate, it is apparent that the photometric amplitudes of and are not tightly coupled – is not detected in the final section of TESS photometry while retains a significant amplitude. While Levenhagen et al. (2011) suggested that may be a circumstellar frequency where LPVs are caused by orbiting material, the fact that this signal is detected in TESS in the absence of any emission (e.g. throughout the entirety of sectors 66+67 prior to the onset of mass ejection) suggests that at least the photometric signal at this frequency is not purely circumstellar.
@@
It should be noted that signals detected in line profile variations may not necessarily correspond to readily detectable photometric signals (and vice-versa). Furthermore, the frequency reported by Levenhagen et al. (2011) is remarkably similar to the frequency we derive for the circumstellar V/R oscillations, and seems to be marginally above the noise in the Fourier spectrum of the pre-whitened light curve for sectors 13 and 66+67 (i.e. with the and signals from Levenhagen and Leister, 2006, removed). It is somewhat difficult to see how an orbital or rotational timescale for the disk can also be present throughout entire sectors of TESS data with an outburst only being present over a few days of one of these sectors.
@@
All of the observed periodic signals in TESS are presumed to be photospheric in origin. The last 3 days of sector 67 (where the disk was starting to build up) were omitted from the frequency analysis, so that all signals from sectors 66+67 in Fig. 4 were recorded during a disk-less state. There was very weak emission present during sector 93, which had dissipated to be no longer detectable throughout sector 94. That is, no mass ejections occurred during these last two sectors. We lack any spectroscopic coverage of TESS sector 13, but there is no evidence of mass ejection in the sector 13 light curve. That is, the end of sector 67 has a sharp rise in flux in Fig. 4, but this is not present in the sector 13 flux.
@@
IV Line Profile Variations Behavior Before, During, and After the Outburst
@@
With two pulsation frequencies and one difference frequency from TESS ( = 0.163 d-1, = 1.644 d-1, = 1.485 d-1), we phased line‑profile maps to examine the behavior of Pav before, during, and after the outburst. In Fig. 1, we see a slow, d motion in the residual line profiles, but a better demonstration of the line profile variations seen in He I 4921 before, during, and after the outburst is in Fig. 5 for the three prominent signals found in the previous section.
@@
Each subplot in Fig. 5 has a combination of the TESS light curve and scaled EW values to show the outburst for He I 4921, with times indicated in gray that coincide with the spectral data used. Then, we include a residual plot for the spectral line as a function of phase calculated with the three prominent TESS frequencies. Lastly, each gray scale residual of the spectral lines is compared to the phased TESS light curve from sectors 66 and 67 when our spectra were obtained.
@@
The strongest photometric signal was d-1, with an amplitude of about 20 ppt from the Fourier analysis of the entire light curve (left column of Fig. 5). The spectroscopic signal appears clearly in each spectral line, with a general trend of an absorption excess starting on the blue side of the line and progressing to the red over the 6.13 d period. The behavior is the same in all three subsets of the data, taken before, during, and after the disk outburst. These figures do contain gaps in phase, which is not unusual for such a small frequency, but this does make the pattern within the LPVs harder to see. One thing to note in the He I and H I lines is that the pulsational behavior extends beyond the literature value of km s-1, out to approximately km s-1, marked with vertical dashed lines. There may be additional line broadening mechanisms besides rotation, and/or it is possible that the value from Zorec et al. (2016) is underestimated. During the outburst, activity and variability can be seen beyond the lines at km s-1, indicating that there must be some disk material present that impacts the observed LPVs as they should not extend to super-rotational velocities.
@@
The second ( d-1) and third ( d-1) highest amplitude frequencies in TESS are more challenging to recover in our spectroscopic data, especially when analyzing sub-sections of the dataset. This is likely due to a number of factors. First, the slow d-1 signal dominates the periodic LPVs. Even in the He I lines analyzed here, the net changes in the EW from circumstellar material have a relatively high amplitude compared to more subtle pulsational LPVs. The signal d-1 is about one d-1 greater than the d-1 signal, which likely causes problems with aliasing. Finally, the fast LPVs discussed in Sec. V.2 are also relatively strong and contribute heavily to the ‘noise’ when searching for LPV patterns at and .
@@
After sub-dividing the spectroscopic data, the only hint of is seen in the pre-outburst data for He I 4921 (top-middle panel, Fig. 5). Although marginally detected, the LPV pattern associated with is similar to that reported in Levenhagen et al. (2011, their Fig. 5), with the alternating light/dark patterns moving across the line profile from blue to red twice per phase. Due at least in part to the issues mentioned in the previous paragraph, is not recovered in the data strings during and after the outburst, and is not apparent in any data sub-strings.
@@
Our observing campaign allowed us to obtain some very dense nights of spectroscopic data throughout the observational campaign. Thus we are also able to see if there are any short timescale variations over individual nights before, during, and after the outburst, although the spectroscopic data are not complete enough in time coverage for determination of periods. Fig. 6 details the H, He I 4713, He I 6678, and Si III 5739 lines. Of the nine nights of dense spectroscopic coverage, we see fast-moving substructures in every time-series, likely related to high-order pulsations in the star that do not have a photometric period due to vanishingly smaller amplitudes from cancellation effects across the stellar disk when the azimuthal order, m, gets larger. As these are seen in every time-series, there does not seem to be a correlation between these line profile variations and the disk-building or dissipation. However, especially in the photospheric Si III line, we see that there are fast-moving prograde (blue-to-red) and retrograde (red-to-blue) features that occur at the same time in this star. These features are seen in Fig. 6 as absorption components that traverse the line profile on the time-scales of hours. The width of the subfeatures is line dependent due to the mass of the atoms making up each line.
V Discussion
@@
V.1 Low-frequency stellar variability
@@
Besides the single outburst, the most conspicuous variability is associated with a 6.13 d period ( d-1). This signal appears in TESS photometry as a sinusoidal modulation of brightness with an amplitude of about 15 – 20 ppt (Figs. 2, 4), and is present in all available TESS sectors (spanning about six years between sector 13 and 94). This same signal is also clearly seen in spectroscopic line profile variations, as first noticed by Levenhagen et al. (2011), and as shown in the dynamical spectra in Fig. 5.
@@
This 0.163 d-1 signal roughly corresponds to the difference between two frequencies observed by TESS, at 1.644 d-1 and 1.485 d-1 (differing by 2%). Such “difference frequencies” are routinely observed in Be stars, especially with the recent proliferation of high-precision space photometry (e.g., Baade et al., 2016, 2017, 2018b). In general the amplitude of difference frequencies is often found to vary significantly over time. This is not the case for Pav – the difference frequency is remarkably regular over the so-far six-year TESS baseline (Fig. 4), and possibly even much longer considering the spectroscopic detection of this signal in data from 1999 to 2001 (Levenhagen et al., 2011), and the photometric detection in Hipparcos data collected in the early 1990s at a frequency of 0.16179 d-1 and an amplitude of 0.016 mag (approximately 16 ppt; Koen and Eyer, 2002). The high observed amplitude of the low-frequency signal relative to all other periodic signals does not conflict with the interpretation that it is a difference frequency. Combination frequencies (including difference frequencies) can easily have higher photometric amplitudes than their parent signals due to differences in geometry/surface patterns (Kurtz et al., 2015).
@@
The line profile variations phased to the period associated with the difference frequency appear to move in a prograde fashion, from blue to red (Figs. 1, 5). From the stellar parameters given in Zorec et al. (2016), the rotation frequency of Pav is 0.46 0.15 d-1, and the critical rotation frequency is 0.62 0.14 d-1 (Paper I). Since the frequency observed at 0.163 d-1 is obviously lower than the rotation frequency, it must be retrograde (unless it is zonal, i.e. , but this seems inconsistent with the LPV pattern). The pattern that moves across the line profile is relatively simple, appearing similar to an 1 or 2 mode, as are regularly seen in Be stars (Rivinius et al., 2003), but at a much lower observed frequency. This low frequency is outside of the range observed in pulsating near-main-sequence B-type stars and the model predictions for them, but as mentioned previously, this is a difference frequency in this star. What is unusual for Pav is that the difference frequency appears to be extremely stable.
@@
Another prominent example of a Be star with a difference frequency is HD 6226. In this case, the difference frequency is not observed directly (either in photometry or line profile variations), but rather is presumed because of two closely spaced pulsation frequencies detected in photometry separated by about 0.0115 d-1 (the corresponding period is 87 d). In HD 6226, outbursts occur with an 87 d period (although some events are skipped or this is modulated on two frequencies). This suggests that the timing of the outbursts is dictated by the difference frequency. In Pav, the difference frequency is over an order of magnitude faster than in HD 6226. Further, mass ejections clearly do not occur every 6 days in Pav. The situation in Pav invokes the idea of “hierarchical clocks” as discussed in Baade et al. (2018b, a), where some unknown mechanism (possibly related to very low difference frequencies) ‘opens a valve’ to allow for mass ejection, where the difference frequency then is able to pump out material for some period of time. Here, the amplitude, which is much higher than those of single-mode stellar pulsations, could be increased by continuous pumping of matter above the photosphere to allow for the disk building to occur. In HD 6226, this hypothetical valve is typically open, while in Pav it is not. One interesting hypothesis that we do not have the ability to test with the current data is that the 0.163 d-1 signal increases in its photometric amplitude at the onset of an outburst as may be seen in Fig. 2. However, the TESS data only cover half of one period of this, so it is impossible to gauge, but similar changes are observed with 25 Ori (B1Vne; Baade et al., 2018b).
@@
The LPV pattern of the slow 0.163 d-1 signal seems qualitatively different from that of the faster 1.644 d-1 signal, most clearly seen in Fig. 5 of (Levenhagen et al., 2011), but also in the right panels of Fig. 5. For the 0.163 d-1 signal, the dark/light patches travel across the line from phase 0 to 1, while for the 1.644 d-1 signal the dark patch (the most evident feature in Levenhagen et al. 2011) and light patch (the most evident feature in the middle column of Fig. 5 in the pre-outburst data) travel across the line profile over just half of the pulsational phase.
@@
It is worth emphasizing one major shortcoming of the TESS photometry and the spectroscopic data analyzed in this work and in Levenhagen et al. (2011). The frequency resolution, which is dictated by the time baseline, is poor. There may therefore be multiple closely spaced frequencies that we detect as only one signal, especially considering that rapid rotation can lead to very small separations between the frequencies associated with consecutive radial orders (e.g., Pápics et al., 2017; Szewczuk et al., 2022). This would also naturally lead to variable amplitudes and/or frequencies in the signals measured by TESS, as unresolved signals beat against each other.
@@
That the LPVs and photometric variations from this difference frequency are so clearly observed and are apparently stable for long periods of time makes Pav an excellent laboratory to study the non-linear mode coupling that seems responsible for the difference frequency (Van Beeck et al., 2024). We are not aware of any other Be star in which the spectroscopic signature of a difference frequency is so evident. A more detailed and systematic investigation of the line profile and brightness variations associated with the 0.163 d-1 signal as well as the two likely ‘parent’ modes (at 1.644 d-1 and 1.485 d-1) may help to answer several important question: Is the pulsation geometry the same for all three observed signals? Why is the difference frequency so regular compared to the majority of Be stars? How do the observed variations map to features on the stellar surface? Given the early spectral type (B2Ve), which is relatively hot compared to the blue edge of the SPB instability strip, what mechanisms excite the two parent modes?
V.2 High-frequency stellar variability